# Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes

### v2

@@ Abstract
Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and out-of-sample predictions. The spatial structure and geometry of the data significantly impact valid statistical inference. Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes. Conversely, data fusion models, while effective, rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets. In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model class utilizing the Hausdorff distance to capture spatial dependence in both point and areal data. We introduce a valid correlation function within the HGP framework, accommodating diverse modeling techniques, including geostatistical and areal models. Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing change of support and data fusion challenges. We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion. The results suggest that the HGP is competitive with specialized models regarding goodness-of-fit and prediction performances. In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
Keywords: Areal data, CAR models, Geostatistics, Spatial modeling, Spatial prediction
@@ 1 Introduction
Accounting for spatial dependence is pivotal in spatial data analysis. Valid statistical inference are deeply influenced by the spatial structure and geometry of the data. Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie, 1993). For areal data, the conditional autoregressive model (CAR) (Besag, 1974) and its variants (Besag et al., 1991; Rodrigues and Assunçao, 2012; Datta et al., 2019; Cruz-Reyes et al., 2023) are often preferred. The nature of the observed spatial data significantly influences the choice of statistical methodology. This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map. This issue, known as change of support (Cressie, 1993, pg. 284–289), requires tailored methods depending on the spatial units involved (Gelfand et al., 2001). Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young, 2002). Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al., 2017).
@@
Both traditional areal models and data fusion models for spatial data exhibit inherent limitations. Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski, 2009; Cruz-Reyes et al., 2023). An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al., 2019). Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al., 2001). Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman, 2018).
@@
We propose a unified approach for analyzing spatial data encompassing areal, point-referenced, and mixed spatial supports. This methodology combines Gaussian processes (GPs) with the Hausdorff distance (Shephard and Webster, 1965). The Hausdorff distance, defined on a reference metric space, effectively measures between the distance between two sets while accounting for their shape and size. For point pairs, it reduces to the distance metric associated to the reference metric space. GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al., 1998), for spatially varying regression coefficients (Gelfand et al., 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand, 2002). The Hausdorff–Gaussian process (HGP) that we propose is a GP with a correlation function defined on the Hausdorff distance between sample units, allowing integrating their shape and size into the analysis. With some ensurance of positiveness definiteness of the resulting covariance matrix, the HGP with a powered exponential correlation can be used in place of GPs to adeptly model areal, point-referenced, and mixed spatial data types, offering a new avenue in analyzing spatial and spatiotemporal data with mixed supports.
@@ 2.1 Hausdorff Distance
The Hausdorff distance is a metric measuring the distance between two sets of points. Specifically, it is the greatest distance from a point in one set to the closest point in the other set. Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines the distance between any two elements $x,y\in D$. For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$. The directed Hausdorff distance from a set $A$ to a set $B$ is $\vec{h}(A,B)=\sup_{a\in A}d(a,B).$ This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances: $h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$ Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$.
In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty, bounded, and compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Sendov, 2004). While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics. Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes.
@@
To demonstrate the efficacy of the Hausdorff distance in measuring distances between various spatial units, Figure 1 presents three simple examples. In each sub-figure, set $A$ is represented as a circle with a radius of 1. Figure 1(a) introduces a set $B$, a circle tangent to $A$. Figure 1(b) examines the distance between two sets having a non-empty intersection. Figure 1(c) features $B$ as an interior point of $A$. The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects. This contrasts with other simpler distance functions between non-empty sets, such as the infimum distance between two sets, denoted $\inf_{x\in X}\inf_{y\in Y}d(x,y)$. This distance does not fulfill the metric property of positiveness condition. In other words, $d(x,y)=0$ does not imply $x=y$. Indeed, employing such distance latter would yield a distance of zero between the three pairs of sets depicted in Figure 1.
@@ 2.2 Hausdorff–Gaussian Process
Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer, 2021). With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object. A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types. Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel, 2006).
@@ 2.3 Correlation Function
The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP as it defines the spatial dependence structure across different locations. In particular, we focus on isotropic correlation functions. Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic. A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie, 1993, Ch. 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$. The latter ensures that spatial correlation diminishes with increasing distance. It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie, 1993, Ch. 2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero, 2006; Li et al., 2023). This raises the question of whether existing frameworks can be adapted or if entirely new formulations are needed to maintain the mathematical rigor in non-Euclidean settings.
@@
$r(h;\varphi,\nu)=\exp\left\{-\frac{h^{\nu}}{\varphi^{\nu}}\right\},$
where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence. To enhance interpretability, we reparametrize this function with $\rho={\log(10)}^{1/\nu}\varphi$. Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range. The PEC function is appealing because the following Proposition 2.1 guarantees its positive definiteness for $\nu\in(1/2,1)$. The proof of the proposition is based on probabilistic embedding of Hausdorff metrics (Backurs and Sidiropoulos, 2016).
Proposition 2.1.
Let $h=h(\mathbf{s},\mathbf{s}^{\prime})$ be the Hausdorff distance between two spatial units, denoted $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$, where $D\subset\rm I\!R^{2}$. The powered exponential correlation function $\exp\{-h^{\nu}/\varphi^{\nu}\}$ is positive definite for $\nu\in(1/2,1)$.
In addition to the theoretical result, an empirical assessment was used to provide evidence for the positive definiteness of this function for an expanded range of $\nu\in(0,1)$. The positive definiteness is satisfied for bigger $\nu$ values when the dependence range $\rho$ is smaller. See details in Section LABEL:sec:sm:emp of the appendix.
@@ 3.2 Bayesian Inferences
Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification. For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$. In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman, 2006). For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$. The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$. In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang, 2004). Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference. The spatial dependence parameter $\rho$ is also inherently difficult. Thus, to avoid oversmoothing, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$. The resulting prior concentrates its mass on lower values $\rho$.
@@
The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7). As the posterior distribution is intractable, we use the No-U-Turn (Hoffman and Gelman, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team, 2023) to draw samples from it. The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation. Additionally, it eliminates the need for hand-tuning, making it a highly convenient sampler. We initialize the parameters with samples from their respective prior distributions. The latent random effects are initialized from a standard Gaussian distribution. The number of samples and the warm-up period of the MCMC algorithm are application dependent. We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al., 2021). Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
@@
The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al., 2017, LOOIC), and the widely applicable information criterion (Watanabe, 2010, WAIC). For both of them, lower values indicate a better fit. We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath, 2016, Ch. 7.4), which would require computationally intensive numerical integration of random effects in some of our models. When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$. Consequently, a negative $\Delta_{G}$ suggests that the HGP model is a better fit for the data. The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
@@
Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al., 2001). The HGP prior for spatial random effects in GLMMs offers a powerful solution. For a set of $m$ unobserved locations $\mathcal{S}^{\ast}=\{\mathbf{s}^{\ast}_{1},\ldots,\mathbf{s}^{\ast}_{m}\}$, we define an $m$-dimensional variable $\mathbf{Y}^{\ast}$ representing the response variable at these locations. Assuming covariates from Equation (4) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
@@
where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters. Details on sampling from this distribution are provided in Section LABEL:sec:sm:hgp-pred of the appendix.

### v3

# Belo Horizonte, MG, Brazil

@@ Abstract
Accurate modeling of spatial dependence is pivotal in analyzing spatial data, influencing parameter estimation and predictions. The spatial structure of the data significantly impacts valid statistical inference. Existing models for areal data often rely on adjacency matrices, struggling to differentiate between polygons of varying sizes and shapes. Conversely, data fusion models rely on computationally intensive numerical integrals, presenting challenges for moderately large datasets. In response to these issues, we propose the Hausdorff-Gaussian process (HGP), a versatile model utilizing the Hausdorff distance to capture spatial dependence in both point and areal data. Integration into generalized linear mixed-effects models enhances its applicability, particularly in addressing data fusion challenges. We validate our approach through a comprehensive simulation study and application to two real-world scenarios: one involving areal data and another demonstrating its effectiveness in data fusion. The results suggest that the HGP is competitive with specialized models regarding goodness-of-fit and prediction performances. In summary, the HGP offers a flexible and robust solution for modeling spatial data of various types and shapes, with potential applications spanning fields such as public health and climate science.
Keywords: CAR models; Geostatistics; Spatial data fusion; Spatial modeling; Spatial prediction
@@ 1 Introduction
Accounting for spatial dependence is pivotal in spatial data analysis. Valid statistical inference are deeply influenced by the spatial structure and geometry of the data. Point-referenced data are typically analyzed using geostatistics methods (e.g., Cressie,, 1993). For areal data, the conditional autoregressive model (Besag,, 1974, CAR) and its variants (Besag et al.,, 1991; Rodrigues and Assunçao,, 2012; Datta et al.,, 2019; Cruz-Reyes et al.,, 2023) are often preferred. The nature of the observed spatial data significantly influences the choice of statistical methodology. This becomes increasingly critical in complex inferential problems, such as predicting a variable in a target map using data from a source map. This issue, known as change of support (Cressie,, 1993, pg. 284–289), requires tailored methods depending on the spatial units involved (Gelfand et al.,, 2001). Spatial data from diverse sources often differ in spatial units or supports, resulting in spatial misalignment (Gotway and Young,, 2002). Data fusion is necessitated when a spatial variable is provided by multiple, potentially misaligned, sources (Moraga et al.,, 2017).
@@
Both traditional areal models and data fusion models for spatial data exhibit inherent limitations. Traditional areal models face challenges with predictions for overlapping spatial units (Bakar and Jin,, 2020), do not account for region size or shape of the areal units and lack interpretable spatial dependence parameters (Assunção and Krainski,, 2009; Cruz-Reyes et al.,, 2023). An exception of the last limitation is the directed acyclic graph auto-regressive (DAGAR) model, which offers an interpretable parameter (Datta et al.,, 2019). Data fusion models, which view areal data as averages of a continuous process, may be unsuitable for discrete outcomes (Gelfand et al.,, 2001). Additionally, their reliance on arbitrary discretization for likelihood evaluation lacks consensus on optimal resolution (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021) and can introduce unquantifiable biases (Gonçalves and Gamerman,, 2018).
@@
We introduce a unified and flexible approach for analyzing spatial data of varying types, encompassing areal, point-referenced, and mixed spatial supports. Our key objective is to develop a model capable of seamlessly handling these diverse data formats within a single framework. This is achieved by combining Gaussian processes (GPs) with the Hausdorff distance (Arutyunov et al.,, 2016). The Hausdorff distance, defined on a reference metric space, effectively measures the distance between two sets while accounting for their shape. For point pairs, it reduces to the distance metric associated to the reference metric space. GPs, recognized for their flexibility, are widely utilized in spatial data modeling, notably as priors in generalized mixed effects models (Diggle et al.,, 1998), in spatially varying regression coefficients (Gelfand et al.,, 2003), and in addressing spatial misalignment issues (Banerjee and Gelfand,, 2002). We propose the Hausdorff-Gaussian process (HGP), a GP with a correlation function defined on the Hausdorff distance between spatial units. This allows us to directly incorporate shape and size information into the analysis. A crucial aspect of our work is to assess the competitiveness of the HGP, particularly when compared to specialized models designed for areal and mixed spatial data. We investigate whether the HGP, using a powered exponential correlation, can effectively model areal and mixed spatial data, thereby offering a novel and versatile tool for analyzing spatial and spatiotemporal data with mixed supports.
@@ 2.1 Hausdorff Distance
The Hausdorff distance is a metric measuring the distance between two sets. Specifically, it is the greatest distance from a point in one set to the closest point in the other set. Consider a metric space $(D,d)$, where $D$ is a spatial region of interest and $d(x,y)$ defines the distance between any two elements $x,y\in D$. For a single point $x$ and a set $A$, define $d(x,A)=\inf_{a\in A}d(x,a)$ as the closest distance from $x$ to any element in $A$. The directed Hausdorff distance from a set $A$ to a set $B$ is
$\vec{h}(A,B)=\sup_{a\in A}d(a,B).$
This measure is not necessarily symmetric. The symmetric Hausdorff distance is defined as the greater of the two directed Hausdorff distances:
$h(A,B)=\max\left\{\vec{h}(A,B),\vec{h}(B,A)\right\}.$
Note that if $A$ and $B$ are both singletons, then $h(A,B)=d(A,B)$. Interestingly, if $A$ and $B$ are compact, there exist $a\in A$ and $b\in B$ such that $h(A,B)=d(a,b)$ (Arutyunov et al.,, 2016).
In spatial analysis, the Hausdorff distance provides a robust and suitable metric over non-empty compact subsets of $D$, denoted as $\mathcal{B}(D)$ (Molchanov,, 2005, Appendix C). While not functioning as a conventional distance in unbounded sets, it is highly effective in bounded contexts, which are prevalent in spatial statistics. Typically, the study region $D$ in spatial statistics is a bounded subset of $\mathbb{R}^{2}$, where the Hausdorff distance excels in quantifying distances between spatial units by capturing both their shapes and sizes. The metric space induced by the tuple $(\mathcal{B}(D),h)$ inherits properties like compactness and completeness from $(D,d)$. Further details regarding the topology induced by the Hausdorff distance can be found in Appendix C inMolchanov, (2005).
@@
To demonstrate the efficacy of the Hausdorff distance in measuring distances between various spatial units, Figure 1 presents three simple examples. In each sub-figure, set $A$ is represented as a circle with a radius of 1. Figure 1a introduces a set $B$, a circle tangent to $A$. Figure 1b examines the distance between two sets having a non-empty intersection. Figure 1c features $B$ as an interior point of $A$. The dashed lines in the figures delineate the Hausdorff distances between the sets, highlighting the metric’s ability to consider the shape, size, and orientation of the objects. This contrasts with other simpler distance functions between non-empty sets, such as the infimum distance between two sets, denoted $\inf_{x\in X}\inf_{y\in Y}d(x,y)$. This distance does not fulfill the metric property of positiveness condition. In other words, $d(x,y)=0$ does not imply $x=y$. Indeed, employing such distance latter would yield a distance of zero between the three pairs of sets depicted in Figure 1.
@@ 2.2 Hausdorff–Gaussian Process
Equation (2) allows a distinction between point-referenced and areal data, a need in data fusion applications (Wang and Furrer,, 2021). With $\mathcal{A}(\cdot)$ being the function that returns the area of a spatial object. A covariate $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$ enables different variances between these data types. Certainly, care is necessary in specifying $v(\mathbf{s})$ to ensure a valid GP (Palacios and Steel,, 2006).
@@ 2.3 Correlation Function
The correlation function $r(\cdot)$ needs to be carefully chosen to ensure the validity of the HGP as it defines the spatial dependence structure across different locations. In particular, we focus on isotropic correlation functions. Therefore, unless explicitly mentioned, all subsequent references to correlation functions will assume they are isotropic. A valid correlation function must be positive definite and meet two key criteria: $r(0)=1$ and $\lim_{h\to\infty}r(h)=0$ (Cressie,, 1993, Ch. 2.3.2), where $h$ is the Hausdorff distance between a pair $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{B}(D)$. The latter ensures that spatial correlation diminishes with increasing distance. It is noteworthy that while most geostatistical correlation functions are constructed for Euclidean spaces (Cressie,, 1993, Ch. 2.5.1), their positive definiteness may not extend to non-Euclidean spaces (Curriero,, 2006; Li et al.,, 2023).
@@
$r(h;\varphi,\nu)=\exp\left\{-\left(\frac{h}{\varphi}\right)^{\nu}\right\},$
where $\nu$ controls the smoothness and $\varphi$ governs the range of the spatial dependence. To enhance interpretability, we reparametrize this function with $\rho={\log(10)}^{1/\nu}\varphi$. Consequently, $\rho$ is the distance at which the spatial correlation between a pair of spatial units drops to $0.10$, commonly referred to as the practical range. The PEC function is appealing because, for $\nu\in(0,1]$, it is positive definite as long as the distance function is conditionally negative definite (Schoenberg,, 1938). To the best of our knowledge, there exist no theoretical results on positive definite functions for the Hausdorff distance. Therefore, we recommend an empirical assessment of the PEC function positive definiteness before analyzing the data.
To this end, we empirically assess the positive definiteness of the PEC function for $\nu\in(0,1]$. We first calculate Hausdorff distance matrices for the areal and fused data to be analyzed in Sections 5.1 and 5.2, respectively. Evaluating the PEC function across these matrices with $\nu\in[0.01,1]$ and $\rho\in[d_{\min},d_{\max}]$ (where $d_{\min}$ and $d_{\max}$ are the dataset’s minimum and maximum distances) yields corresponding correlation matrices. We then compute each matrix’s smallest eigenvalue. A negative eigenvalue suggests the corresponding PEC function is not positive definite. However, even valid covariance functions can be ill-conditioned or nearly rank-deficient at dense locations or when there exists a severely strong spatial dependence (Banerjee et al.,, 2012).
Figure 2: Smallest eigenvalue of the powered exponential correlation matrix under different combinations of smoothness ($\nu$) and practical range ($\rho$) for areal and fused data. The transparency indicates regions where $\lambda_{1}<0$.
Figure 2 displays the smallest eigenvalues ($\lambda_{1}$) of the PEC-induced correlation matrices for both datasets. The fused data exhibits positive eigenvalues for all $(\nu,\rho)$ pairs. Conversely, the areal data reveals some small negative eigenvalues (range 0.0099-0.0099 to 0.00006-0.00006) under severely strong spatial dependence and at $\nu\approx 1$, the border of the parametric space. Therefore, our empirical study suggests the positive definiteness of the PEC function for $\nu\in(0,1]$.
@@ 3.2 Bayesian Inferences
Prior distributions for all the model parameters are to be specified to complete the Bayesian model specification. For fixed effects regression coefficients ($\bm{\beta}$), we employ uncorrelated zero-mean Normal priors with marginal variances $10^{2}$. In the homoscedastic case, on $\sigma$ we place a half-t prior with $3$ degrees of freedom, denoted $t_{+}(3)$ (Gelman,, 2006). For the hereroscedastic model (2), we set independent standard normal priors for $\alpha_{0}$ and $\alpha_{1}$. The PEC dependence is encoded through $\bm{\delta}=\{\rho,\nu\}$. In geostatistics, smoothness parameters for correlation functions are typically fixed, and their estimation properties are rarely reported (Zhang,, 2004). Due to this and the scope of this paper, we specify separate models for different values of $\nu$ ensuring a positive definite PEC and select the best fit for inference. The spatial dependence parameter $\rho$ is also inherently difficult. Thus, to avoid oversmoothing, we employ an exponential prior for $\rho$ with mean $\lambda_{\rho}=-\rho_{0}/\log(p_{\rho})$, where $\rho_{0}$ is an upper-bound set to four-fifths of the largest observed Hausdorff distance, and $p_{\rho}$ is the probability that $\rho$ exceeds $\rho_{0}$. The resulting prior concentrates its mass on lower values $\rho$.
@@
The estimation of the spatial GLMM parameters is obtained from the posterior distribution defined in Equation (7). As the posterior distribution is intractable, we use the No-U-Turn (Homan and Gelman,, 2014) Markov Chain Monte Carlo (MCMC) sampler available in Stan (Team,, 2023) to draw samples from it. The No-U-Turn algorithm samples all the model parameters jointly and efficiently exploits the parameter space using automatic differentiation. Additionally, it eliminates the need for hand-tuning, making it a highly convenient sampler. We initialize the parameters with samples from their respective prior distributions. The latent random effects are initialized from a standard Gaussian distribution. The number of samples and the warm-up period of the MCMC algorithm are application dependent. We assess the convergence of the chains using the split-$\hat{R}$ diagnostic (Vehtari et al.,, 2021). Finally, the parameters’ point estimates and 95% credible intervals (CI) are obtained as, respectively, the median and percentiles ($2.5$ and $97.5$, unless otherwise stated) of the marginal MCMC samples.
@@
The assessment of goodness-of-fit (GoF) in the simulation study and applications involved two model comparison criteria: the leave-one-out information criterion (Vehtari et al.,, 2017, LOOIC), and the widely applicable information criterion (Watanabe and Opper,, 2010, WAIC). For both of them, lower values indicate a better fit. We did not use the deviance information criterion due to its assumption of a small number of parameters relative to the sample size (McElreath,, 2016, Ch. 7.4), which would require computationally intensive numerical integration of random effects in some of our models. When comparing a single candidate model to the HGP, we utilize the difference in the GoF metric $G$ between the HGP and the given model, denoted by $\Delta_{G}$. Consequently, a negative $\Delta_{G}$ suggests that the HGP model is a better fit for the data. The competing methods depend on the type of application we are working with and will be mentioned in the upcoming sections.
@@
Predicting spatial phenomena accurately at unobserved locations is crucial in many applications, especially in change of support settings (Gelfand et al.,, 2001). The HGP prior for spatial random effects in GLMMs offers a powerful solution. For a set of $m$ unobserved locations $\mathcal{S}^{\ast}=\{\mathbf{s}^{\ast}_{1},\ldots,\mathbf{s}^{\ast}_{m}\}$, we define an $m$-dimensional variable $\mathbf{Y}^{\ast}$ representing the response variable at these locations. Assuming covariates from Equation (4) are available, we can generate predictions accounting for uncertainty using the posterior predictive distribution of $\mathbf{Y}^{\ast}$:
@@
where $\mathbf{z}^{\ast}={(z(\mathbf{s}^{\ast}_{1}),\ldots,z(\mathbf{s}^{\ast}_{m}))}^{\top}$ is a realization of the spatial random effect at the new locations, and $\bm{\theta}$ represents the model parameters. Details on sampling from this distribution are provided in Section S.1 of the Appendix Material.